# On the exact self-similar finite-time blowup of the Hou–Luo model with smooth profiles

### v2

@@ 3.2. Estimates of $b(f)$ and $c(f)$
Recall that $\rho(x)=(1+|x|)^{1+\delta_{\rho}}$. Denote $\delta=\|f_{1}-f_{2}\|_{L^{\infty}_{\rho}}=\|\rho(f_{1}-f_{2})\|_{L^{\infty}}$. Since $(1+x^{2}/2)^{-2}\leq f_{i}\leq 1$, $i=1,2$, we have
$|f_{1}(x)-f_{2}(x)|\leq\min\left\{x^{2}\,,\,\delta(1+|x|)^{-1-\delta_{\rho}}\right\}$
Then,
@@ 3.3. Monotonicity and convexity
Next, we derive the key property of $\bm{R}$, that it preserves the monotonicity in $x$ and convexity in $x^{2}$ on $[0,+\infty)$ for functions in $\mathbb{D}$. This nice property of $\bm{R}$ essentially arises from a kernel analysis of the linear map $\bm{T}$ as shown in the following lemma, which is a restatement of its counterpart in [9]. We still provide the proof here for the sake of completeness.
@@
Given $f\in\mathbb{D}$, $\bm{T}(f)^{\prime}(x)\leq 0$ on $[0,+\infty)$, and $\bm{T}(f)(\sqrt{s})$ is convex in $s$.
@@
[…]
@@
$\frac{g(x)-1}{x^{2}}=\frac{g(x)-g(0)}{x^{2}}\leq\lim_{x\rightarrow+\infty}\frac{g^{\prime}(x)}{2x}=\frac{1}{2},$
@@
For any $f\in\mathbb{D}$, $\bm{M}(f)^{\prime}(x)\leq 0$ on $[0,+\infty)$, and $\bm{M}(f)(\sqrt{s})$ is convex in $s$.
@@
For any $f\in\mathbb{D}$, $\bm{R}(f)^{\prime}(x)\leq 0$ on $[0,+\infty)$, and $\bm{R}(f)(\sqrt{s})$ is convex in $s$. Moreover, $\bm{R}(f)(x)\geq\bm{M}(f)(x)$ for all $x$.
@@
[…]
@@
Note that, by Corollaries 3.6 and 3.7, the function $dm(x)/(g(x)+d-1)$ is non-increasing in $x$. Hence,
@@
$\tilde{A}(x)=A(x)-B(x)m(x),\quad\tilde{B}(x)=x^{2}B(x).$ (3.15)
@@
From the above we also find that
@@
By the concavity of $g(\sqrt{s})$ in $s$ (Corollary 3.6), $g^{\prime}(x)/x$ and $(g(x)-1)/x^{2}$ are both non-increasing in $x$. Moreover, the function $m(x)/(3g(x)+d-1)$ is also non-increasing in $x$. Hence, $\tilde{A}^{\prime}(x)/\tilde{B}^{\prime}(x)$ is non-increasing in $x$. This again implies that $\tilde{A}(x)/\tilde{B}(x)\geq\tilde{A}^{\prime}(x)/\tilde{B}^{\prime}(x)$, and thus, for $x>0$,
@@
Write $g=\bm{G}(f)$. We first upper bound $\bm{T}(f)^{\prime}(x)$ in two ways using some estimates in [9] (more precisely, in the proof of [9, Lemma 3.6]). We restate the detailed calculations below for the reader’s convenience. On the one hand, we can use the calculations in the proof of Lemma 3.5 to get that, for $x>0$,
@@
$-\bm{T}(f)^{\prime}(x)\geq\left(\frac{1}{3\sqrt{2}\pi x^{3}}\right)^{1/5}(1-f(x))^{1/2}\cdot\left(\frac{x|f^{\prime}(x)|}{2\pi}\right)^{4/5}=\frac{1}{\pi}\left(\frac{x}{48\sqrt{2}}\right)^{1/5}|f^{\prime}(x)|^{4/5}(1-f(x))^{1/2}.$
@@
We have used the $x$-dependent upper bound of $c(f)$ in Lemma 3.3. Plugging in $x=1$ and using $f^{\prime}(1)\leq-\eta$, we get
@@ 3.4. Decay estimates
Let $r=\bm{R}(f)$, $m=\bm{M}(f)$, $g=\bm{G}(f)$, and $d=d(f)$. Bear in mind that $g(x)$ is non-decreasing on $[0,+\infty)$, and $1=g(0)\leq g(x)\leq g(+\infty)=2d$. Define
@@
We have used the fact $a=27\eta/8<1/7$ so that $43(1+a)/(1-a)<$ 4. ∎
@@
The first inequality above owes to the non-increasing property of $f$ on $[0,+\infty)$, and the second inequality follows from $m_{1}(x)\leq f(x)\leq 5x^{-\delta_{0}}$. Note that for $0\leq y\leq t_{0}x$,
@@
[…]
@@ 3.6. Existence of a fixed point
$-\frac{f^{\prime}(x)}{2x}\leq\frac{f(0)-f(x)}{x^{2}}\leq\min\{1\ ,\frac{1}{x^{2}}\},\quad x>0.$
@@
Furthermore, since $|f_{n}^{\prime}(x)|\leq 2$ on $[0,X_{m}]$, we can apply Ascoli’s theorem to select a sub-sequence $\{f_{n_{m,k}}\}_{k=1}^{+\infty}$ of $\{f_{n_{m-1,k}}\}_{k=1}^{+\infty}$ such that $\|\rho(f_{n_{m,i}}-f_{n_{m,j}})\|_{L^{\infty}}\leq 2\epsilon_{m}$ for any $i,j\geq 1$. Then the diagonal sub-sequence $\{f_{n_{m,m}}\}_{m=1}^{+\infty}$ is a Cauchy sequence in the $L_{\rho}^{\infty}$-norm. This proves that $\mathbb{D}$ is sequentially compact. ∎
@@
By Theorem 3.18 and Lemma 3.19, the nonempty set $\mathbb{D}$ is convex, closed and compact in the $L_{\rho}^{\infty}$-norm, and $\bm{R}$ continuously maps $\mathbb{D}$ into itself. The Schauder fixed-point theorem then guarantees that $\bm{R}$ has a fixed point in $\mathbb{D}$. ∎
4. Properties of the fixed-point solution
Throughout this section, we always denote by $f$ a fixed point of $\bm{R}$ in $\mathbb{D}$, i.e. $f=\bm{R}(f)$. We will use this fixed-point relation to derive finer characterizations of $f$. As usual, we write $g=\bm{G}(f),m=\bm{M}(f)$ and $d=d(f)$. Provided that $f=\bm{R}(f)$, we have
@@ 4.1. Regularity
Also note that $m(x)/g(x)\leq m(0)/g(0)=1$. We immediately have $m^{\prime}\in H^{p}(\mathbb{R})$. Rearranging the first equation of (4.1), we obtain
@@ 4.2. Asymptotic behavior
We have used the fact that the non-negative function $\frac{1}{t^{\delta}}\ln\left|\frac{t+1}{t-1}\right|$ is integrable on $[0,+\infty)$ for any $\delta\in(0,2)$. This proves the first result.
@@
Since the limit $D_{\delta}:=\lim_{x\rightarrow+\infty}x^{1+\delta}f(x)$ exists and is finite, we know $C_{\delta}:=\sup_{x\in\mathbb{R}}\ x^{1+\delta}f(x)<+\infty$ and $\lim_{x\rightarrow+\infty}(tx)^{1+\delta}f(tx)=D_{\delta}$ for any fixed $t>0$. Also note that the function $t^{-\delta}\ln\left|\frac{t+1}{t-1}\right|$ is absolutely integrable on $[0,+\infty)$ for any $\delta\in(0,2)$. We then use the dominated convergence theorem to obtain
@@
We can now compute the decay rates of $\bm{M}(f)$ and $\bm{R}(f)$ based on the asymptotic behavior of $\bm{G}(f)$.
@@
Next, we prove the asymptotic behavior of $r$. Let $A(x),B(x)$ be defined as in (3.13). Recall that $A(x)$ is non-decreasing on $[0,+\infty)$, and thus $A(+\infty)$ is well-defined. We then use (4.5) to find that
[…]
@@ 4.3. Estimates of $d(f)$
Multiply both sides of the equation above by $2/\pi$ and then integrating them over $[0,+\infty)$ yields
@@
$-\frac{c_{\omega}}{c_{l}}=\frac{b(f)-c(f)}{2(b(f)+c(f))}=\frac{d(f)-1}{2d(f)}=\delta_{d}.$
@@ 5. Numerical verification
More precisely, each iteration is computed in the order of (2.9). After the scheme converges numerically, the self-similar profiles $\omega,v$ and the scaling factors $c_{l},c_{\omega}$ are then recovered as in Proposition 3.1 and renormalized as in (2.4). Note that a similar fixed-point method for computing the self-similar profiles of the 1D gCLM model was proposed in [9] by the same authors.
@@
The second purpose is to check whether the self-similar profiles determined by the fixed-point method in this paper and those obtained by solving the dynamic rescaling equations as in [3] are identical under proper rescaling. It turns out that the numerical profiles obtained by the iterative method (5.1) and those obtained by solving the dynamic rescaling equations (data stored in [2]) under consistent normalization ($f(0)=m(0)=1$) are almost identical up to only mesh-point-wise errors at the level of $10^{-5}$. The errors are likely due to the differences in the discretization methods. This convincingly supports our conjecture.
We remark that the most numerically expensive step in one iteration is to compute the function $\bm{T}(f^{(n)})$, which involves the evaluation of a linear transform with a dense kernel at all mesh points. The mesh points are distributed adaptively over a sufficiently large one-sided interval $[0,10^{16}]$ (solutions are truncated to 00 for $x\geq 10^{16}$). An analogous computation (recovering $u$ from $\omega$) is also needed in every time step when solving the dynamic rescaling equations (2.3) numerically, and it takes more than tens of thousands of time steps for the solution to converge in time. Hence, our method is empirically more efficient than numerically solving the dynamic rescaling equations in obtaining approximate self-similar profiles.
Finally, we provide some plots of the numerically constructed profiles to verify and visualize some of their theoretically proved properties. Figure 5.1 plots the numerically obtained fixed point $f$ and the corresponding $\bm{M}(f)$ in $x$ and in $s=x^{2}$ respectively, verifying that they are both monotone decreasing in $x$, convex in $x^{2}$, and lower bounded by $(1+x^{2}/2)^{-2}$ for $x\geq 0$. Figure 5.1 plots the corresponding $\bm{G}(f)$ in a similar way, verifying that it is monotone increasing in $x$, concave in $x^{2}$, and upper bounded by $21+x^{2}/$ 2 for $x\geq 0$. Figure 5.3 demonstrates asymptotic decay rates of $f$ and $\bm{M}(f)$ for sufficiently large $x$, verifying the statements in Theorem 4.4.
@@
Figure 5.2. $\bm{G}(f)$ of the numerically obtained fixed point $f$ plotted (a) in coordinate $x$ and (b) in coordinate $s=x^{2}$. The dashed line represents the upper bound $21+x^{2}/2=1+s/$ 2.

### v3

@@ 3.2. Estimates of $b(f)$ and $c(f)$
Recall that $\rho(x)=(1+|x|)^{1+\delta_{\rho}}$. Denote $\delta=\|f_{1}-f_{2}\|_{L^{\infty}_{\rho}}=\|\rho(f_{1}-f_{2})\|_{L^{\infty}}$. Since
$\frac{1}{(1+x^{2}/2)^{2}}\leq f_{i}(x)\leq 1,\quad i=1,2,$
we have
$|f_{1}(x)-f_{2}(x)|\leq\min\left\{x^{2}\,,\,\delta(1+|x|)^{-1-\delta_{\rho}}\right\}.$
Hence,
@@ 3.3. Monotonicity and convexity
Next, we derive the key property of $\bm{R}$, that it preserves the monotonicity in $x$ and convexity in $x^{2}$ on $0,+\infty)$ for functions in $\mathbb{D}$. This nice property of $\bm{R}$ essentially arises from a kernel analysis of the linear map $\bm{T}$ as shown in the following lemma, which is a restatement of its counterpart in [16]. We still provide the proof here for the sake of completeness.
@@
Given $f\in\mathbb{D}$, $\bm{T}(f)^{\prime}(x)\leq 0$ on $0,+\infty)$, and $\bm{T}(f)(\sqrt{s})$ is convex in $s$.
@@
For any $f\in\mathbb{D}$, $\bm{G}(f)^{\prime}(x)\geq 0$ on $0,+\infty)$, and $\bm{G}(f)(\sqrt{s})$ is concave in $s$. As a consequence, $1\leq\bm{G}(f)(x)\leq\min\{1+x^{2}/2\,,\,2d(f)\}$ for all $x$.
@@
$\frac{g(x)-1}{x^{2}}=\frac{g(x)-g(0)}{x^{2}}\leq\lim_{x\rightarrow 0}\frac{g^{\prime}(x)}{2x}=\frac{1}{2},$
@@
For any $f\in\mathbb{D}$, $\bm{M}(f)^{\prime}(x)\leq 0$ on $0,+\infty)$, and $\bm{M}(f)(\sqrt{s})$ is convex in $s$.
@@
For any $f\in\mathbb{D}$, $\bm{R}(f)^{\prime}(x)\leq 0$ on $0,+\infty)$, and $\bm{R}(f)(\sqrt{s})$ is convex in $s$. Moreover, $\bm{R}(f)(x)\geq\bm{M}(f)(x)$ for all $x$.
@@
$\begin{split}A(x)&:=\int_{0}^{x}dy^{d-1}\cdot\frac{m(y)}{g(y)}\exp\left(\big{(}d-1\big{)}\int_{0}^{y}\frac{1-g(z)}{zg(z)}\,\mathrm{d}{z}\right)\,\mathrm{d}{y},\\ B(x)&:=x^{d}\exp\left(\big{(}d-1\big{)}\int_{0}^{x}\frac{1-g(y)}{yg(y)}\,\mathrm{d}{y}\right),\end{split}$ (3.13)
@@
Note that, by Corollaries 3.6 and 3.7, the function dm(x)/(g(x)+d1)dm(x)/(g(x)+d-1) is non-increasing in $x$. Hence,
@@
$\tilde{A}(x):=A(x)-B(x)m(x),\quad\tilde{B}(x):=x^{2}B(x).$ (3.15)
@@
This concludes that $\bm{R}(f)(x)\geq\bm{M}(f)(x)$ for all $x$. From the above we also find that
@@
By the concavity of $g(\sqrt{s})$ in $s$ (Corollary 3.6), $g^{\prime}(x)/x$ and $(g(x)-1)/x^{2}$ are both non-increasing in $x$. Moreover, the function m(x)/(3g(x)+d1)m(x)/(3g(x)+d-1) is also non-increasing in $x$. Hence, $\tilde{A}^{\prime}(x)/\tilde{B}^{\prime}(x)$ is non-increasing in $x$. This again implies that $\tilde{A}(x)/\tilde{B}(x)\geq\tilde{A}^{\prime}(x)/\tilde{B}^{\prime}(x)$, and thus, for $x>0$,
@@
Write $g=\bm{G}(f)$. We first upper bound $\bm{T}(f)^{\prime}(x)$ in two ways using some estimates in [16] (more precisely, in the proof of [16, Lemma 3.6]). We restate the detailed calculations below for the reader’s convenience. On the one hand, we can use the calculations in the proof of Lemma 3.5 to get that, for $x>0$,
@@
$\bm{T}(f)^{\prime}(x)\geq\left(\frac{1}{3\sqrt{2}\pi x^{3}}\right)^{1/5}(1-f(x))^{1/2}\cdot\left(\frac{x|f^{\prime}(x)|}{2\pi}\right)^{4/5}=\frac{1}{\pi}\left(\frac{x}{48\sqrt{2}}\right)^{1/5}|f^{\prime}(x)|^{4/5}(1-f(x))^{1/2}.$
@@
We have used the upper bound of $c(f)$ in Lemma 3.3. Plugging in $x=1$ and using $f^{\prime}(1)\leq-\eta$, we get
@@ 3.4. Decay estimates
Let $r=\bm{R}(f)$, $m=\bm{M}(f)$, $g=\bm{G}(f)$, and $d=d(f)$. Bear in mind that $g(x)$ is non-decreasing on $0,+\infty)$, and $1=g(0)\leq g(x)\leq g(+\infty)=2d$. Define
@@
We have used the fact $a=27\eta/8<1/7$ so that 3(1+a)/(1a)<43(1+a)/(1-a)<4. ∎
@@
The first inequality above owes to the non-increasing property of $f$ on $0,+\infty)$, and the second inequality follows from $m_{1}(x)\leq f(x)\leq 5x^{-\delta_{0}}$. Note that for $0\leq y\leq t_{0}x$,
@@
The claim that $\bm{R}$ maps $\mathbb{D}$ into itself follows from Corollaries 3.8, 3.14, 3.16, 3.10, and 3.12. In order to prove $\mathbb{D}$ is nonempty, it suffices to show that $m_{1}\in\mathbb{D}$. Recall the definition (3.1) of $m_{1}$. Firstly, it is not hard to check that the function $g_{1}$ satisfies $g_{1}(x)$ is non-decreasing on $0,\infty)$ and $g_{1}(\sqrt{s})$ is concave in $s$, which implies $m_{1}(x)$ is non-increasing on $0,\infty)$ and $m_{1}(\sqrt{s})$ is convex in $s$ (by a similar argument as in the proof of Corollary 3.7). Secondly, a direct derivative calculation shows that $m_{1}^{\prime}(1)\leq-\eta$ (for that $\eta$ is very small). Thirdly, through the proofs of preceding lemmas and corollaries, it is straightforward to check that $m_{1}(x)\leq\min\big{\{}(1+3/\delta_{1})(|x|/L_{1})^{-1-\delta_{1}}\,,\,5|x|^{-\delta_{0}}\big{\}}$. The above together imply that $m_{1}\in\mathbb{D}$, and the theorem is thus proved. ∎
@@ 3.6. Existence of a fixed point
$\frac{f^{\prime}(x)}{2x}\leq\frac{f(0)-f(x)}{x^{2}}\leq\min\{1\ ,\frac{1}{x^{2}}\},\quad x>0.$
@@
Furthermore, since $|f_{n}^{\prime}(x)|\leq 2$ on $0,X_{m}]$, we can apply Ascoli’s theorem to select a sub-sequence $\{f_{n_{m,k}}\}_{k=1}^{+\infty}$ of $\{f_{n_{m-1,k}}\}_{k=1}^{+\infty}$ such that $\|\rho(f_{n_{m,i}}-f_{n_{m,j}})\|_{L^{\infty}}\leq 2\epsilon_{m}$ for any $i,j\geq 1$. Then the diagonal sub-sequence $\{f_{n_{m,m}}\}_{m=1}^{+\infty}$ is a Cauchy sequence in the $L_{\rho}^{\infty}$-norm. This proves that $\mathbb{D}$ is sequentially compact. ∎
@@
By Theorem 3.18 and Lemma 3.19, the nonempty set $\mathbb{D}$ is convex, closed and compact in the $L_{\rho}^{\infty}$-norm, and $\bm{R}$ continuously maps $\mathbb{D}$ into itself. The Schauder fixed-point theorem (Fact B.1) then guarantees that $\bm{R}$ has a fixed point in $\mathbb{D}$. ∎
4. Properties of a fixed-point solution
In this section, we derive finer properties such as smoothness and far-field behavior of a fixed point $f=\bm{R}(f)$ in $\mathbb{D}$ by using the fixed-point relation. We remark that all properties to be established below apply to any possible fixed point of $\bm{R}$ in $\mathbb{D}$, though the uniqueness of a fixed point of $\bm{R}$ is conjectured.
As usual, we write $g=\bm{G}(f),m=\bm{M}(f)$ and $d=d(f)$. Provided that $f=\bm{R}(f)$, we have
@@ 4.1. Regularity
Also recall that $g(x)\geq g(0)=1$ and $m(x)\leq m(0)=1$. Moreover, it is not hard to check that $\xi^{\prime}\in H^{p}(\mathbb{R})$ implies $(\xi-\xi(0))/x\in H^{p}(\mathbb{R})$ for any suitable function $\xi$. We then immediately have $m^{\prime}\in H^{p}(\mathbb{R})$. Rearranging the first equation of (4.1), we obtain
@@ 4.2. Asymptotic behavior
We have used the fact that the non-negative function $\frac{1}{t^{\delta}}\ln\left|\frac{t+1}{t-1}\right|$ is integrable on $0,+\infty)$ for any $\delta\in(0,2)$. This proves the first result.
@@
Since the limit $D_{\delta}:=\lim_{x\rightarrow+\infty}x^{1+\delta}f(x)$ exists and is finite, we know $C_{\delta}:=\sup_{x\in\mathbb{R}}\ x^{1+\delta}f(x)<+\infty$ and $\lim_{x\rightarrow+\infty}(tx)^{1+\delta}f(tx)=D_{\delta}$ for any fixed $t>0$. Also note that the function $t^{-\delta}\ln\left|\frac{t+1}{t-1}\right|$ is absolutely integrable on $0,+\infty)$ for any $\delta\in(0,2)$. We then use the dominated convergence theorem to obtain
@@
We can now compute the decay rates of $\bm{M}(f)$ and $\bm{R}(f)$ by studying the asymptotic behavior of $\bm{G}(f)$.
@@
Next, we prove the asymptotic behavior of $r$. Let $A(x),B(x)$ be defined as in (3.13). Recall that $A(x)$ is non-decreasing on $0,+\infty)$, and thus $A(+\infty)$ is well-defined. We then use (4.5) to find that
$A(+\infty)$ $=\int_{0}^{+\infty}dy^{d-1}\cdot\frac{1}{g(y)^{2}}\exp\left(\big{(}d+1\big{)}\int_{0}^{y}\frac{1-g(z)}{zg(z)}\,\mathrm{d}{z}\right)\,\mathrm{d}{y}$
@@ 4.3. Estimates of $d(f)$
Multiply both sides of the equation above by $2/\pi$ and then integrating them over $0,+\infty)$ yields
@@
$\frac{c_{\omega}}{c_{l}}=\frac{b(f)-c(f)}{2(b(f)+c(f))}=\frac{d(f)-1}{2d(f)}=\delta_{d}.$
@@ 5. Numerical verification
More precisely, each iteration is computed in the order of (2.9). After the scheme converges numerically, the self-similar profiles $\omega,v$ and the scaling factors $c_{l},c_{\omega}$ can be recovered as in Proposition 3.1 and renormalized as in (2.4). Note that a similar fixed-point method for computing the self-similar profiles of the 1D gCLM model was proposed in [16] by the same authors.
@@
The second purpose is to check whether the self-similar profiles determined by the fixed-point method in this paper and those obtained by solving the dynamic rescaling equations (2.3) as in [7] are identical under proper rescaling. For the profiles constructed in [7], we obtain the profile data, denoted by $(\bar{\omega},\bar{v},\bar{u},\bar{c}_{l},\bar{c}_{\omega})$, from their open source code [5]. We then construct the corresponding functions $\bar{f}=\bar{\omega}/x$, $\bar{m}=2\bar{v}/(\bar{c}_{l}x)$, and $\bar{g}=(\bar{c}_{l}+\bar{u}/x)/(c_{l}+\bar{u}^{\prime}(0))$ (as in (2.6)) and properly rescale them (according to (2.4)) so that $\bar{f}(0)=1$ and $\lim_{x\rightarrow 0}\bar{g}^{\prime}(x)/x=\bar{g}^{\prime\prime}(0)=1$. Finally, we compare $(\bar{f},\bar{m})$ with our numerically obtained fixed-point solution $(f,m)$ where $m=\bm{M}(f)$ (see Figure 5.1). It turns out that the numerical profiles obtained by the iterative method (5.1) and those obtained by solving the dynamic rescaling equations under consistent normalization are almost identical only up to grid-point errors at the level of $10^{-7}$. The errors are likely due to the differences in the discretization methods. This convincingly supports our conjecture on uniqueness. Moreover, the ratio $c_{l}/|c_{\omega}|$ numerically computed by our fixed-point iteration is approximately $2.99869$ (rounding to five decimal places), which is very close to the estimated value $\bar{c}_{l}/|\bar{c}_{\omega}|=2.99870$ given in [7]. In fact, our result satisfies the rigorous computer-assisted estimate $|(c_{l}/|c_{\omega}|)-2.99870|\leq 6\times 10^{-5}$ established in [7].
(a) $f(x)$ versus $\bar{f}(x)$
(b) $f(x)-\bar{f}(x)$
(c) $m(x)$ versus $\bar{m}(x)$
(d) $m(x)-\bar{m}(x)$
Figure 5.1. Comparison between $(f,m)$ and $(\bar{f},\bar{m})$. Here, $f$ and $m=\bm{M}(f)$ are obtained numerically using the fixed-point iteration scheme (5.1); $\bar{f}$ and $\bar{m}$ are constructed from the data in [5] under proper normalization. The maximal difference between $f$ and $\bar{f}$ (or between $m$ and $\bar{m}$) over all grid points (including those beyond the plotting range in these figures) is below $10^{-6}.$
We remark that the most numerically expensive step in one iteration is to compute the function $\bm{T}(f^{(n)})$, which involves the evaluation of a linear transform with a dense kernel at all mesh points. The mesh points are distributed adaptively over a sufficiently large one-sided interval $0,10^{16}]$ (solutions are truncated to 0 for $x\geq 10^{16}$). An analogous computation (recovering $u$ from $\omega$) is also needed in every time step when solving the dynamic rescaling equations (2.3) numerically, and it takes more than tens of thousands of time steps for the solution to converge in time. Hence, our method is empirically more efficient than numerically solving the dynamic rescaling equations in obtaining approximate self-similar profiles.
Finally, we provide some plots of the numerically constructed profiles to verify and visualize some of their theoretically proved properties. Figure 5.2 plots the numerically obtained fixed point $f$ and the corresponding $\bm{M}(f)$ in coordinates $x$ and $s=x^{2}$ respectively, verifying that they are both monotone decreasing in $x$, convex in $s=x^{2}$, and lower bounded by $(1+x^{2}/2)^{-2}$ for $x\geq 0$. Figure 5.3 plots the corresponding $\bm{G}(f)$ in a similar way, verifying that it is monotone increasing in $x$, concave in $x^{2}$, and upper bounded by $1+x^{2}/2$ for $x\geq 0$. Figure 5.4 demonstrates asymptotic decay rates of $f$ and $\bm{M}(f)$ for sufficiently large $x$, verifying the statements in Theorem 4.4.
@@
Figure 5.3. $\bm{G}(f)$ of the numerically obtained fixed point $f$ plotted (a) in coordinate $x$ and (b) in coordinate $s=x^{2}$. The dashed line represents the upper bound $1+x^{2}/2=1+s/2$.